Strong base is dissolved in 765 mL of 0.400 M weak acid (Ka 4.23x 10) to make a buffer with a pH of 4.19. Assume that the volume remains constant when the base is added. HA(aq) + OH-(aq) → H2O()-A-(aq) Calculate the pKa value of the acid and d les of acid initially present. Number Number pK, = mol HA When the reaction is complete, what is the concentration ratio of conjugate base to acid? Number HA How many moles of strong base were initially added? Number mol OH
Answer #1

pKa = - log Ka = - log 4.23 \times 10^{-5} = 4.37
Number of moles of acid HA present initially = 0.400 \times \dfrac {765}{1000} = 0.306 moles
pH = pKa + log \dfrac {[A^-]}{[HA]} \\ 4.19 = 4.37 + log\dfrac {[A^-]}{[HA]} \\ \dfrac {[A^-]}{[HA]} = 0.655
Let x moles of strong base were initially added
\dfrac {[A^-]}{[HA]} = 0.655 = \dfrac {x}{0.306-x} \\ 1.526x = 0.306-x \\ x = 0.121
Hence, 0.121 moles of strong base were added initially.